OpenStudy (anonymous):

t=1/8x-3/4y, solve for x

OpenStudy (anonymous):

t + 3/4y = 1/8x -> x = 8t + 6y

OpenStudy (anonymous):

Could you please explain it further, I'm not understanding where the ">" came from.

OpenStudy (anonymous):

oops, sorry, that was just an arrow saying that the next step follow from this, ok, again, add the 3/4y to both sides, and you get t + 3/4y = 1/8x, now multiply both sides with 8, and you get 8t + (8 * 3/4)y = x

OpenStudy (anonymous):

and in (8 * 3/4) you can cancel the 8 with the 4 and you get (2 * 3)

OpenStudy (anonymous):

Ohh Ok thank you.

OpenStudy (anonymous):

It's been a pleasure :P